If it lands on a six 140 times then the estimated probability of a six is 140/400 = 0.35
Why you prefer Normal Distribution over the other distributions in Statistics?
For theoretical reasons (such as the central limit theorem), any variable that is the sum of a large number of independent factors is likely to be normally distributed. For this reason, the normal distribution is used throughout statistics, natural science, and social science as a simple model for complex phenomena.
Sir names 1st Senior second Junior third the third when is it ever the second?
I am a 2nd and not a junior. I have the same name as my Grandfather. If I were to have the same name as my father I would be a junior. If I give my son the same as myself he would be a junior and a 3rd.
What is the chance of being shot by a random gunman?
Pretty slim in the US or most of Europe and Asia. If you are in Africa or the Middle East you would have a significantly higher risk. Still relatively small though, you're much more likely to die in a car acccident than from a random gunman.
The probability of observing a z value equal to or more extreme than 1.50 is 0.1336
Why are random samples more representative of the population than other kinds of samples?
A 'random' sample - covers all age groups, genders, and other criteria. A 'targeted' sample might only cover a small part of the population.
What is the probability of drawing 2 aces from a 52 card deck?
Let A= getting an ace the first time
and B= getting an ace the second
We are looking to for the probaliity of getting A and B that is P(A and B)
We know P(A and B) = P(A) . P(B|A)
= (4/52) . (3/51) = 1/122 = .00452
NOTE that P(B|A) is the conditional probability of getting an ace the second time given that you got an ace the first time.
Your risk of getting pregnant is the same as any other time when you are fertile. Unless you are menopausal or pregnant, you have a chance of becoming pregnant from having sex.
What is the probability of drawing a red number card from a deck of cards?
13 out of 52.
_____________________
It should be 20 out of 52 if Ace is included as a number card.
Not including ace the probability is 18 out of 52.
What is the probability of rolling a 5 or 6 on a die?
1 out of six for both five and six well that's for the original die wich is 1 of each number from 1 to 6 it depends how many sides it has
What is the probability of drawing a king or a 9 from a standard 52 card deck of cards?
The probability of drawing a king or a nine from a standard deck of 52 cards is (4 + 4) in 52, or 8 in 52, or 2 in 13, or about 0.1538.
The answer is 4! (4 factorial), the same as 4x3x2x1, which equals 24 combinations.
The answer is 24 and this is how:
A b c d
A b d c
A c d b
A c b d
A d c b
A d b c
B c d a
B c a d
B d a c
B d c a
B a c d
B a d c
C d a b
C d b a
C a b d
C a d b
C b d a
C b a d
D a b c
D a c b
D b c a
D b a c
D c a b
D c b a
Assuming that the chosen are over the age of 85, the probability is
(0.28)8 = 0.000038
Probability is a number in the range [0, 1]. The question gives a probability (240) which is way outside this range and so is not valid.
99 percent confidence interval Population mean 24.4 to 38.0 find the mean sample?
if the confidence interval is 24.4 to 38.0 than the average is the exact middle:
31.2, and the margin of error is 6.8
Find the next number in the following sequence 2-5-10-17-26?
37 because they are increased by consecutive odd numbers each time.
What is one standard deviation above the mean 700?
The mean alone is not enough to provide an answer.
Is a good way to increase a confidence interval is to increase the sample size?
No, the opposite is true.